# WikiAutoGen: Towards Multi-Modal Wikipedia-Style Article Generation

### v1

@@ 4.5 Multi-modal Article Writing
After generating the textual content, the next step is to identify and incorporate relevant images to enhance the article’s readability and expressiveness. This multimodal integration process involves several stages: image positioning proposal, image retrieval, image selection, and finally, an article refinement step that seamlessly integrates the images and text.
Image positioning proposal. After obtaining the complete textual article, the next step is to utilize an LLM-based agent to suggest appropriate placements and corresponding descriptions for relevant images. These initial proposals are then refined through interaction with the multi-perspective self-reflection module, which evaluates their relevance, coherence, and engagement from the reader’s perspective.
Image retrieval. We then retrieve relevant images by performing searches based on multiple sources, including general image search engines, Wikipedia, and also the websites mentioned in the references. After that, we can obtain a list of relevant images.
Image selection. We first use the CLIP model [32] to retrieve candidate images by computing their similarity to the query caption, selecting the top-3 most relevant images. Subsequently, we leverage a multi-modal model [16] to further evaluate these candidates and select the most suitable image for inclusion in the article.
Paper polish. After finalizing images selection, we integrate the chosen images into the article and proceed to a polishing stage. Due to potential discrepancies between textual content and visual figures, we employ a multi-modal model to revise the entire article, enhancing coherence and consistency across modalities. Additionally, this multimodal model interacts with the multi-perspective self-reflection module, obtaining editorial feedback to further refine the integrated content.
@@ 5.1 Experiment setup
Implementation Details. For the language model (LM) components of WikiAutoGen, we employ zero-shot prompting implemented using the DSPy framework [19] in conjunction with GPT-4o [16], GPT-4o-mini and GPT-o3-mini [28]. Specifically, we use GPT-o3-mini for the multi-perspective self-reflection module due to its strong reasoning capabilities, GPT-4o for the multimodal knowledge exploration tasks, and GPT-4o-mini for all other remaining tasks. WikiAutoGen retrieves real-time web information using Serper’s API [^1] https://serper.dev/, with each query returning up to 5 web pages. Throughout the experiments, we maintain consistent settings by fixing the LM temperature at 1.0 and the $top\_p$ value at 0.9.
Evaluation metrics. We evaluate the multimodal articles by separately assessing their textual and visual content.
@@
- Image quality evaluation. As there are no existing benchmarks specifically designed for evaluating multimodal Wikipedia-style article generation, we propose an evaluation method inspired by the textual evaluation frameworks [33, 17] to evluate image quality. Specifically, we assess image quality based on four criteria: image-text coherence, engagement, helpfulness, and information supplement (the image’s ability to provide additional useful information beyond the textual context).
Baselines We compare our approach with four representative LLM-based baseline frameworks for automated expository writing on our WikiSeek benchmark:
@@
Since these baselines originally lack multimodal capabilities, we equip them with image retrieval functionalities to enable a fair multimodal comparison. Specifically, each baseline can also retrieve image via Google image search, extract relevant metadata, and search for images based on generated textual descriptions.
@@ 5.2 Experimental results
Textual content comparison on WikiSeek. We demonstrate our results in the Table 2. The results indicate that our approach consistently achieves the highest average scores across all evaluation dimensions—content quality, informativeness, reliability, and engagement—highlighting its comprehensive effectiveness in multimodal article generation.
For text-only inputs, our method achieves an average score of 78.73, significantly outperforming the best baseline (Co-Storm, 69.96) by approximately 8.8 points, demonstrating superior coherence, alignment, and informativeness. In image-only scenarios, the improvement is even more pronounced, with our model obtaining 77.49, surpassing the strongest baseline (oRAG, 48.38) by approximately 29.1 points, reflecting its exceptional capability to extract meaningful textual insights from visual content. For combined image-text topics, our model maintains its advantage with an average score of 78.82, showing a clear improvement (15.05 points) over the next-best baseline (Co-Storm, 63.77). Overall, the substantial performance gains indicate that our multimodal approach excels at synthesizing cohesive and engaging content from diverse inputs.
Image content comparison on WikiSeek. Table 4 compares the performance of image content across different article generation methods on our WikiSeek benchmark. Our WikiAutoGen consistently achieves the highest scores across all evaluation metrics for all three input modalities.
Specifically, our method significantly improves upon baseline methods, outperforming the next-best method by approximately 11.34 points on image-text coherence (72.24 vs. Storm’s 59.67) under the image-text topics. Similarly, for image-only topics, our approach excels particularly in helpfulness (77.63) and coherence (71.90), demonstrating the model’s superior capability in selecting images that meaningfully complement textual content and provide additional useful information. Overall, these results underscore our method’s effectiveness in seamlessly integrating images to enhance multimodal coherence, engagement, and informativeness, substantially advancing the quality and reader experience of multimodal articles.
@@ 5.3 Ablation studies
Figure 3: Ablation study across different data difficulty levels. We compare our WikiAutoGen with Storm and OmniThink on three difficulty categories: hard (300–500 characters), very hard (100–300 characters), and extremely hard ($<$100 characters).
Ablation on different components of WikiAutoGen. We conduct an ablation study to analyze the individual contributions of three key components on textual article generation from image-only topics. The components include multi-agent knowledge exploration, outline proposal, and multi-perspective self-reflection. Results are shown in Table 3. Specifically, without the outline proposal, the system simply retrieves a single image and extracts its description from metadata. Without multi-agent knowledge exploration, only a single fixed agent responds to the asker.
Introducing the multi-agent module significantly improves average performance from 48.78 (no modules enabled) to 66.56 (+17.78 points), highlighting its effectiveness in collaborative knowledge exploration. Using the outline proposal alone boosts performance further to 72.88 (+24.10 points), underscoring its importance for content structuring. The self-reflection module individually achieves 71.60 (+22.82 points), indicating its strength in refining coherence and consistency. Combining all three modules results in the highest performance (77.49), demonstrating their complementary roles in generating coherent, informative, and engaging articles from image-based inputs.
Ablation on different difficulty levels. In our WikiSeek benchmark, topics are grouped into three difficulty levels based on article length (character count): hard (300–500 characters), very hard (100–300 characters), and extremely hard (less than 100 characters), with 100 examples per level.
We evaluate text-only topics and report the average text evaluation performance in Fig. 3. The results indicate that our WikiAutoGen consistently outperforms baseline methods across all difficulty levels. Notably, as the difficulty increases (from hard to extremely hard), the performance gap between WikiAutoGen and Storm widens from 12.35 points to 15.35 points, and similarly increases compared to other baselines. This demonstrates the robustness and stability of WikiAutoGen in effectively handling highly challenging and underexplored topics.
@@ 5.4 Human evaluation
To assess the quality of the generated Wikipedia-style articles, we conduct a human evaluation study using Amazon Mechanical Turk (AMT) [^2] https://www.mturk.com/. We randomly sample 100 text-based topics from the WikiSeek dataset and perform pairwise comparisons between articles generated by our method (WikiAutoGen), Storm, and OmniThink.
For each topic, three independent participants evaluate the articles in a randomized order. Participants first answer the question: “Do you think adding images would improve comprehension of the topic?” Results shown in Fig. 4 (left) indicate that 99% of participants agree images enhance comprehension.
Additionally, participants answer multiple-choice questions including: (1) Which article is the easiest to understand? (2) Which article is the most engaging in terms of narrative, examples, or overall presentation? (3) Which article provides the most comprehensive background information and in-depth analysis? (4) Which article is your overall favorite? The order of these questions is randomized to reduce bias. As illustrated in Fig. 4 (right), participants consistently prefer articles generated by WikiAutoGen over Storm and OmniThink across all evaluation criteria.
Figure 4: Human evaluation study. We randomly sample text-only topics and conduct comparative evaluations between WikiAutoGen, Storm, and OmniThink. Left: Participants respond to the question, “Do you think adding images would improve comprehension of the topic?”. Right: Participants answer multiple-choice questions evaluating readability, engagement, informativeness, and overall preference.
@@ 6 Conclusion
In this paper, we introduced WikiAutoGen, a comprehensive multi-agent framework designed for automated multi-modal Wikipedia-style article generation. Our approach integrates both visual and textual content, significantly enhancing the depth, informativeness, and engagement of generated articles. To address limitations in existing methods, we proposed a novel multi-perspective self-reflection mechanism, ensuring improved coherence, reliability, and overall article quality. Furthermore, we presented WikiSeek, a challenging multimodal benchmark specifically crafted to evaluate the performance of models in generating content for less-explored topics. Experimental results demonstrated that WikiAutoGen substantially outperforms existing state-of-the-art methods across textual and visual evaluation metrics. Particularly notable were the improvements in content quality, informativeness, and reader engagement, validating the effectiveness of integrating multimodal inputs and iterative self-reflection.

### v2

@@ 4.5 Multi-modal Article Writing
Following textual generation, we incorporate relevant visual content to enhance the article’s readability and expressiveness. This multimodal integration process involves several stages: image positioning proposal, image retrieval, image selection, and finally, an article refinement step that seamlessly integrates the images and text.
Image positioning proposal. After generating the complete textual article, an LLM-based agent is employed to propose appropriate placements and corresponding descriptions for relevant images. These initial proposals are then refined through interaction with the multi-perspective self-reflection module, which evaluates their relevance, coherence, and engagement from the reader’s perspective.
Image retrieval. We then retrieve relevant images by performing searches based on multiple sources, including general image search engines, Wikipedia, and the websites mentioned in the references. After that, we can obtain a list of relevant image candidates.
Image selection. We first use the CLIP model [27] o rank retrieved images based on semantic similarity to the generated captions, selecting the top-3 candidates. Subsequently, we leverage a multi-modal model [13] to further evaluate these candidates and select the most contextually appropriate image for inclusion in the article.
Article Polishing. After finalizing image selection, we integrate the chosen images into the article and proceed to a polishing stage. Due to potential discrepancies between textual content and visual figures, we employ a multi-modal model to revise the entire article, enhancing coherence and consistency across modalities. Additionally, this multimodal model interacts with the multi-perspective self-reflection module, obtaining editorial feedback to further refine the integrated content.
@@ 5.1 Experiment setup
Implementation Details. For the language model (LM) components of WikiAutoGen, we employ zero-shot prompting implemented using the DSPy framework [15] in conjunction with GPT-4o [13], GPT-4o-mini, and GPT-o3-mini [23]. Specifically, we use GPT-o3-mini for the multi-perspective self-reflection module due to its strong reasoning capabilities, GPT-4o for the multimodal knowledge exploration tasks, and GPT-4o-mini for all other remaining tasks. WikiAutoGen retrieves real-time web information via Serper’s API [^1] https://serper.dev/, with each query returning up to 5 web pages. Throughout the experiments, we maintain consistent settings by fixing the LM temperature at 1.0 and the top_ptop\_p value at 0.9. For evaluation, we utilize GPT-4o as the evaluator. To further validate the results, we conduct more evaluations using two distinct evaluators, Gemini2.5-Flash [8] and Prometheus2 [16], as detailed in Appendix B.
Evaluation metrics. We evaluate the generated multimodal articles through separate assessments of their textual and visual content.
@@
- Image quality evaluation. As there are no existing benchmarks specifically designed for evaluating multimodal Wikipedia-style article generation, we propose an evaluation method inspired by the textual evaluation frameworks [28, 14] to evaluate image quality. Specifically, we assess image quality based on four criteria: image-text coherence, engagement, helpfulness, and information supplement (the image's ability to provide additional useful information beyond the textual context).
Baselines We compare WikiAutoGen with four representative LLM-based baseline frameworks for automated expository writing on our WikiSeek benchmark:
@@
Since these baselines originally lack multimodal capabilities, we equip them with image retrieval functionalities to enable a fair multimodal comparison. Specifically, each baseline can also retrieve images via Google image search, extract relevant metadata, and search for images based on generated textual descriptions.
@@ 5.2 Experimental results
Textual content comparison on WikiSeek. We demonstrate our results in the Table 2. The results indicate that WikiAutoGen consistently achieves the highest average scores across all evaluation dimensions—content quality, informativeness, reliability, and engagement—highlighting its comprehensive effectiveness in article generation.
For text-only inputs, our WikiAutoGen achieves an average score of 78.73, significantly outperforming the best baseline (Co-Storm, 69.96) by approximately 8.8 points, demonstrating superior coherence, alignment, and informativeness. In image-only scenarios, the improvement is even more pronounced, with WikiAutoGen obtaining 77.49, surpassing the strongest baseline (oRAG, 48.38) by approximately 29.1 points, reflecting its exceptional capability to extract meaningful textual insights from visual content. For combined image-text topics, our model maintains its advantage with an average score of 78.82, showing a clear improvement (+ 15.05 points) over the next-best baseline (Co-Storm, 63.77). Overall, the substantial performance gains demonstrate that WikiAutoGen excels at synthesizing cohesive and engaging content across diverse inputs.
Image content comparison on WikiSeek. Table 4 compares the performance of image content across different article generation methods. Our WikiAutoGen consistently achieves the highest scores across all image evaluation metrics for all three input modalities.
Specifically, our method significantly improves upon baseline methods, outperforming the next-best method by approximately 11.34 points on image-text coherence (72.24 vs. Storm’s 59.67) under the image-text topics. Similarly, for image-only topics, our approach excels particularly in helpfulness (77.63) and coherence (71.90), demonstrating WikiAutoGen's superior capability in selecting images that meaningfully complement textual content and provide additional useful information. Overall, these results highlight WikiAutoGen's effectiveness in integrating images to enhance coherence, engagement, and informativeness, substantially advancing the overall quality and readability of multimodal articles.
@@ 5.3 Ablation studies
Figure 3: Ablation study across different data difficulty levels. We compare our WikiAutoGen with Storm and OmniThink on three difficulty categories: hard (300–500 characters), very hard (100–300 characters), and extremely hard (<<100 characters).
Ablation on different components of WikiAutoGen. We conduct an ablation study to analyze the individual contributions of three core components to textual article generation from image-only topics. Specifically, we examine the impact of components, including multi-agent knowledge exploration, outline proposal, and multi-perspective self-reflection. Results are shown in Table 3. Specifically, without the outline proposal, the system simply retrieves a single image and extracts its description from metadata. Without multi-agent knowledge exploration, only a single static agent responds to the asker.
Incorporating the multi-agent module improves performance from 48.78 (baseline without modules) to 66.56 (+17.78 points), highlighting its effectiveness in collaborative knowledge exploration. Using the outline proposal alone further increases performance to 72.88 (+24.10 points), underscoring its importance for content structuring. The self-reflection module individually achieves 71.60 (+22.82 points), indicating its strength in refining coherence and consistency. Combining all three modules results in the highest performance (77.49), validating their complementary roles in generating coherent, informative, and engaging articles from image-only inputs.
Ablation on different difficulty levels. In our WikiSeek benchmark, topics are grouped into three difficulty levels based on article length (character count): hard (300–500 characters), very hard (100–300 characters), and extremely hard (fewer than 100 characters), with 100 examples per level. We evaluate text-only inputs and present the average textual evaluation results in Figure 3. The results indicate that our WikiAutoGen consistently outperforms all baseline methods across all three difficulty levels. Notably, as the topic difficulty increases (from hard to extremely hard), the performance gap between WikiAutoGen and Storm widens from 12.35 points to 15.35 points, with similar trends observed against other baselines. These results demonstrate the robustness and stability of WikiAutoGen in effectively handling highly challenging and underexplored topics.
5.4 Compared with Commercial Deep Research
Method Text as Topic Image as Topic Image-Text as Topic Average Compute Time
OpenAI 92.75 91.95 94.50 93.06  $\sim\,30\ \mathrm{min}$
Google 81.91  $\sim\,12\ \mathrm{min}$
Grok 88.35 81.70 86.13 85.06  $\sim\,10\ \mathrm{min}$
WikiAutoGen 88.58 89.55 88.89 89.01  $\sim\,\mathbf{8}\ \mathrm{min}$
Table 5: Comparison of commercial models' article generation performance with Prometheus2 [16].
[…]
@@ 5.5 Human evaluation
To evaluate the quality of the generated Wikipedia-style articles, we conduct a human evaluation study via Amazon Mechanical Turk (AMT) [^2] https://www.mturk.com/. We randomly sample 100 text-only topics from the WikiSeek benchmark dataset and perform pairwise comparisons between articles generated by our method (WikiAutoGen), Storm, and OmniThink. Each topic is evaluated by three independent participants in randomized order. Participants first answer the question: ``Do you think adding images would improve comprehension of the topic?'' As shown in Figure 4 (left), 97.7% of participants agree that images improve topic comprehension.
Additionally, participants answer multiple-choice questions, including: (1) Which article is the easiest to understand? (2) Which article is the most engaging in terms of narrative, examples, or overall presentation? (3) Which article provides the most comprehensive background information and in-depth analysis? (4) Which article is your overall favorite? The question order is randomized to mitigate potential evaluation bias. As illustrated in Figure 4 (right), participants consistently prefer articles generated by WikiAutoGen over those by Storm and OmniThink across all evaluation criteria.
Figure 4: Human evaluation study. We randomly sample text-only topics and conduct comparative evaluations between WikiAutoGen, Storm, and OmniThink. Left: Participants respond to the question, ``Do you think adding images would improve comprehension of the topic?". Right: Participants answer multiple-choice questions evaluating readability, engagement, informativeness, and overall preference.
@@ 6 Conclusion
In this paper, we introduced WikiAutoGen, a comprehensive multi-agent framework designed for automated multimodal Wikipedia-style article generation. WikiAutoGen integrates both visual and textual content, significantly enhancing the depth, informativeness, and engagement of generated articles. To address key limitations in prior work, we proposed a novel multi-perspective self-reflection mechanism, which systematically improves article coherence, reliability, and overall quality. Furthermore, we presented WikiSeek, a challenging multimodal benchmark specifically crafted to evaluate the performance of models in generating content for less-explored topics. Experimental results demonstrated that WikiAutoGen substantially outperforms existing state-of-the-art baselines across both textual and visual evaluation metrics. Particularly notable were the improvements in content quality, informativeness, and reader engagement, validating the effectiveness of integrating multimodal inputs and iterative self-reflection.